### v1

# GLiDRE: Generalist Lightweight model for Document-level Relation Extraction

Abstract
Relation Extraction (RE) is a fundamental task in Natural Language Processing, and its document-level variant poses significant challenges, due to the need to model complex interactions between entities across sentences. Current approaches, largely based on the ATLOP architecture, are commonly evaluated on benchmarks like DocRED and Re-DocRED. However, their performance in zero-shot or few-shot settings remains largely underexplored due to the task’s complexity. Recently, the GLiNER model has shown that a compact NER model can outperform much larger Large Language Models. With a similar motivation, we introduce GLiDRE, a new model for document-level relation extraction that builds on the key ideas of GliNER. We benchmark GLiDRE against state-of-the-art models across various data settings on the Re-DocRED dataset. Our results demonstrate that GLiDRE achieves state-of-the-art performance in few-shot scenarios. Our code is publicly available [^1] https://github.com/robinarmingaud/glidre.
GLiDRE: Generalist Lightweight model for Document-level Relation Extraction
Robin Armingaud, Romaric Besançon Université Paris-Saclay, CEA, List, F-91120, Palaiseau, France {robin.armingaud, romaric.besancon}@cea.fr
1 Introduction
Document-level Relation Extraction (DocRE) is a challenging task in natural language processing that involves identifying relationships between entities across multiple sentences. In contrast, traditional sentence-level relation extraction, evaluated on supervised datasets like TACRED (Zhang et al., 2017) and SemEval-2010 Task 8 (Hendrickx et al., 2010) or few-shot benchmarks such as FewREL (Han et al., 2018) and Wiki-ZSL (Chen and Li, 2021), typically considers a single entity pair per instance, often without negative examples (i.e. sentences where no relation is present). These settings closely resemble relation classification rather than relation extraction. While recent works improve existing benchmarks, like FewRel 2.0 (Gao et al., 2019) which addresses the negative examples issue, they still operate under the constraint of classifying a fixed pair of entities.
DocRE presents a more realistic and complex scenario, especially relevant for real-world applications such as Information Extraction (IE) in biomedical or industrial texts, where relationships often span sentence boundaries and the entity pairs of interest are not pre-identified. The complexity of DocRE is due to the quadratic growth of negative pairs with the number of entities in a document. Evaluation is commonly conducted on datasets like DocRED (Yao et al., 2019) and its revised version Re-DocRED (Tan et al., 2022b), which reduces the issue of false negatives. These two datasets contain more complex documents with an average of 19 entities and approximately 390 candidate pairs per text, for a total of 3,053 annotated Wikipedia articles in the training set.
State-of-the-art approaches to DocRE generally extend the model ATLOP (Zhou et al., 2021), which introduces Localized Context Pooling, using attention mechanisms to enhance relation representations, and an adaptive-thresholding loss that adjusts decision thresholds dynamically based on entity pairs. Subsequent work has extended ATLOP in various directions. For example, DREEAM (Ma et al., 2023) uses evidences to enhance performance, while KD-DocRE (Tan et al., 2022a) uses axial attention to model interdependencies among entity pairs, adopts an adaptive focal loss to address class imbalance, and integrates knowledge distillation.
Figure 1: Model architecture. GLiDRE uses a dual-encoder structure: a text encoder generates token embeddings, while a label encoder computes embeddings for relation types. An entity’s representation is formed by pooling the embeddings of its constituent tokens and mentions. For each entity pair, their representations are concatenated and processed by a feed-forward network before final scoring. The model can optionally incorporate a localized context pooling module, inspired by ATLOP, to enrich the relation representation.
However, if recent works tackle sentence-level relation extraction in low-resource settings (Boylan et al., 2025, Li et al., 2024a, Lan et al., 2023), DocRE in similar settings remains largely unexplored. Although latest LLMs show strong capabilities in zero-shot settings for tasks such as NER and RE (Sainz et al., 2024, Zhou et al., 2023, Wang et al., 2023, Wei et al., 2023), their performance on zero-shot DocRE, especially on the DocRED or Re-DocRED dataset is still limited (Li et al., 2023, Xue et al., 2024). The recent success of GLiNER (Zaratiana et al., 2024) has motivated this work. GLiNER deals with the task of NER as matching entity type representations to textual span representations in the latent space. This approach allows better representations than LLMs due to its use of bidirectional encoders and it solves the scalability issues of autoregressive models, surpassing much larger LLMs in zero shot settings. Moreover, as a zero-shot model, GLiNER is not restricted to a fixed set of relation labels and can generalize to arbitrary relation types at inference time. Based on this work, we introduce GLiDRE, a model inspired by the bi-encoder version of GLiNER [^2] https://blog.knowledgator.com/meet-the-new-zero-shot-ner-architecture-30ffc2cb1ee0, adapted for DocRE. Our contributions in this paper are :
  • We propose GLiDRE, a new and efficient model for document-level relation extraction, inspired by the core principles of GLiNER. We also adapt and test key components from ATLOP to improve relation representations.
  • We conduct extensive experiments on the Re-DocRED benchmark, demonstrating that GLiDRE achieves state-of-the-art results in challenging few-shot scenarios.
  • We demonstrate that smaller, specialized encoder models can offer an efficient alternative to larger LLMs for complex, document-level relation extraction and we present an up-to-date evaluation of these models on the Re-DocRED dataset in a zero-shot setting.
2 Methodology
2.1 Document Level Relation Extraction
Let a document be represented as a sequence of tokens $D$. Within this document, there is a set of $m$ pre-identified entities $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each entity $e_{i}$ is associated with one or more mentions in the text.
The objective of DocRE is to identify valid relational triplets. Each triplet is of the form $(e_{h},r,e_{t})$, where $e_{h},e_{t}\in E$ are the head and tail entities respectively ($h$ can be equal to $t$), and $r$ is the relation that holds between them. Based on this definition, DocRE is as a multi-label classification task.
2.2 GLiDRE architecture
GLiDRE extends the bi-encoder variant of GLiNER. This architecture employs two separate transformers, one for encoding the document and another for encoding relation labels, which offers several benefits : unlike the uni-encoder version, the bi-encoder approach avoids concatenating labels and document tokens into a single input. Instead, we encode the document and labels separately. This eliminates the need for special separation tokens and overcomes encoder context length limits, permitting an effectively unlimited label set. Moreover, label embeddings can be precomputed, accelerating inference. However, the bi-encoder architecture increases memory usage and lacks cross-attention between labels. As a result, it may struggle to disambiguate semantically similar labels.
Label and word representations
Label embeddings are obtained by mean-pooling over the token representations constituting label names. The computed label embeddings are passed through a two-layer feedforward network if the dimension of the label encoder’s output differs from the configured latent space dimension of GLiDRE. For document tokens, we adopt the standard strategy used in GLiNER and other standard NER models : for words split into subwords, we extract the representation of the first subword.
Relation representations
Given contextual embeddings $H\in\mathbb{R}^{L\times D}$ from the document encoder, where $L$ is the sequence length and $D$ is the hidden dimension of the document encoder, we construct a representation for each candidate relation defined by a pair of entities.
First, for each mention of an entity, we compute the mention representation by pooling the embeddings of its constituent words.
If an entity has multiple mentions, its final representation $h_{e}$ is derived by taking the mean of all its mention representations.
Finally, the representation $h_{r}$ for the relation between $e_{h}$ and $e_{t}$, is obtained by concatenating their respective entity representations and passing them through a two-layer feedforward network.
$h_{r}=\text{FFN}(h_{e_{h}}\otimes h_{e_{t}})$ (1)
where $\otimes$ denotes vector concatenation.
Relation Scoring
To determine whether a candidate entity pair $(e_{h},e_{t})$ instantiates a given relation type $t$, we measure the following matching score between the relation representation $h_{r}$ and the embedding of relation type $h_{t}$ :
$s(e_{h},e_{t},t)=\sigma(h_{r}^{\top}h_{t})$ (2)
where $\sigma(x)=(1+e^{-x})^{-1}$ is the sigmoid function. A relation $t$ is assigned to the entity pair $(e_{h},e_{t})$ if
$s(e_{h},e_{t},t)>\tau$ (3)
where $\tau$ is a predefined decision threshold.
2.3 Loss Function
To address the issue of class imbalance within DocRE datasets, we use the Focal Loss function proposed by Lin et al. (2017) during training. The Focal Loss is a dynamically scaled cross-entropy loss designed to down-weight the contribution of well-classified examples and concentrating the training on hard examples. It is defined as:
$\text{FL}(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ (4)
where $p_{t}$ represents the model’s estimated probability for the ground-truth class. The focusing parameter $\gamma$ adjusts the rate at which easy examples are down-weighted and the weighting factor $\alpha_{t}$ balances the importance of positive and negative examples.
2.4 Localized Context Pooling
To create more refined representations, we also explore and adapt methods inspired by ATLOP Zhou et al. (2021) including Localized Context Pooling. Localized Context Pooling uses the attention scores from the encoder to focus on the most relevant parts of the document for a given entity pair.
Let $A\in\mathbb{R}^{L\times L}$ be the document-level attention matrix from the last encoder layer, averaged across all attention heads. For the head entity $e_{h}$ and tail entity $e_{t}$, we extract their corresponding attention vectors $A_{h}$ and $A_{t}$ by averaging the attention scores over their mention tokens. A joint attention vector $\alpha$ is computed via an element-wise product:
$\alpha=A_{h}\odot A_{t}$ (5)
This $\alpha$ is then normalized and used to compute a weighted sum of the contextual embeddings, creating a localized context vector $c_{\text{loc}}$:
$c_{\text{loc}}=\sum_{i=1}^{L}\frac{\alpha_{i}}{\sum_{j=1}^{L}\alpha_{j}}H_{i}$ (6)
Finally, the refined representations for the head $h^{\prime}_{e_{h}}$ and tail $h^{\prime}_{e_{t}}$ entities are created by concatenating their initial representations with the localized context vector and passing them through distinct two-layer feedforward network:
$h^{\prime}_{e_{h}}=\tanh(\text{FFN}_{h}(h_{e_{h}}\otimes c_{\text{loc}}))$ (7)
$h^{\prime}_{e_{t}}=\tanh(\text{FFN}_{t}(h_{e_{t}}\otimes c_{\text{loc}}))$ (8)
The final relation representation $h_{r}$ is then computed using Equation 1.
2.5 Pretraining data
To construct a large-scale pre-training corpus, we use a semi-automated annotation methodology inspired by Zhou et al. (2023). We randomly sample documents from FineWeb Penedo et al. (2024), a high-quality dataset built from filtered and deduplicated English Common Crawl archives.
We prompt the Mistral-Small-24B-Instruct-2501 model to generate annotations for both entities and the relations between them in a structured JSON format (the detailed prompt is provided in appendix A.1). The raw outputs are filtered to remove instances containing malformed JSON or documents that exceed a predefined length threshold.
The resulting dataset contains 136,404 documents with highly diverse labels, featuring 76,497 unique relation types. The most frequent relation types include types such as IS_LOCATED_IN, WORKS_FOR, PART_OF and CONTAINS. We use this dataset to pretrain our model.
3 Experiments
3.1 Dataset
We conduct our experiments on the English benchmark Re-DocRED Tan et al. (2022b), a human-revised version of the DocRED benchmark that addresses its high false-negative rate, logical inconsistencies and coreference errors by re-annotating all 4,053 documents in the original training and evaluation splits. Re-DocRED retains the same relation schema and splits but substantially increases the number of annotated relation triples per document.
3.2 Evaluation Settings
We evaluate the performance of our model across three distinct experimental settings, with a particular focus on low-data regimes where its advantages are most pronounced. First, in a fully supervised setting, we fine-tune on the entire training set and compare against strong baselines. Second, to assess data efficiency and few-shot generalization, we fine-tune on randomly sampled subsets of the training data containing 1, 5, 10, 50, 100, 500, or 1000 examples. Finally, we examine zero-shot performance by comparing against larger LLMs without any task-specific fine-tuning, highlighting the model’s competitiveness in scenarios with little or no labeled data.
Evaluation Metrics
We use the standard metrics from the Re-DocRED dataset, defined in Tan et al. (2022b): F1 refers to the micro-averaged F1 score on the recognition of relation triples, Ign_F1 corresponds to the F1 score when ignoring the triples in the test set that appear also in the training set.
Implementation Details
All experiments are conducted on a single NVIDIA H100 GPU with a batch size of 16. For the pre-training phase, the model is trained for 50,000 steps. For finetuning, we train for 10,000 steps. We use two learning rates $1\times 10^{-5}$ for the model’s encoders and $1\times 10^{-4}$ for all other layers. For the few-shot and fully supervised experiments, we report the average and standard deviation over 5 runs with different seeds. The model checkpoint yielding the best performance on the development set and a decision threshold of 0.5 are used for the evaluation on the test set. We use two English models : DebertaV3 Large He et al. (2021) for document encoder and BGE-Large V1.5 Xiao et al. (2023) for label encoder similar to the bi-encoder version of GLiNER. The final model comprises approximately 800 million parameters. The pre-training phase requires approximately 24 hours, while finetuning on the ReDocRED dataset takes 3.5 hours.
3.3 Results
3.3.1 Fully supervised setting
We evaluate the performance of our proposed model, GLiDRE, by comparing it against several strong baselines on the Re-DocRED dataset under a fully supervised protocol. Specifically, we include:
  • ATLOP (Zhou et al., 2021): Formulates document-level relation extraction as a semantic-segmentation task, introducing localized context pooling to capture entity-focused local contexts and an adaptive thresholding loss to learn dynamic decision thresholds.
  • KD-DocRE (Tan et al., 2022a): Leverages an axial attention module to model inter-entity dependencies across sentences, applies an adaptive focal loss to mitigate class imbalance and uses knowledge distillation to incorporate distantly supervised data.
  • DREEAM (Ma et al., 2023): Guides the model’s attention with evidence sentences as supervisory signals in a memory-efficient manner and employs a self-training strategy to learn evidence retrieval without explicit annotations.
  • TTM-RE (Gao et al., 2024): Introduces a Token Turing Machine memory module that augments document representations with external memory tokens, paired with a noise-robust loss tailored for positive–unlabeled distantly supervised settings.
Additionally, we benchmark against GLiREL (Boylan et al., 2025), a GLiNER sentence-level relation extraction adaptation. It is important to note that GLiREL was originally conceived for sentence-level relation extraction and operates with a shorter context length. For document-level predictions, it relies on aggregating relations based on gold-standard coreference information.
We also benchmark against recent methods that leverage LLMs. Specifically, we compare with the framework introduced by Li et al. (2024b), which involves fine-tuning Llama models (Touvron et al., 2023) using Low-Rank Adaptation (Hu et al., 2022) and enhancing the model’s performance by employing a classifier to exhibit potential relations and guide the fine-tuned LLM.
Table 1: Fully supervised results on the Re-DocRED dataset. We compare GLiDRE with strong DocRE models, LLM-based methods and GLiREL. F1 scores are reported on both the development and test sets. Results are taken from the original papers, except for DocRE models, which are reported following Gao et al. (2024).
Model Dev F1 Dev Ign F1 Test F1 Test Ign F1
DocRE Models
ATLOP (Zhou et al., 2021) $76.15_{\pm 0.23}$ $75.88_{\pm 0.23}$ $77.81_{\pm 0.71}$ $76.13_{\pm 0.28}$
KD-DocRE (Tan et al., 2022a) $77.88_{\pm 0.42}$ $77.12_{\pm 0.49}$ $78.28_{\pm 0.72}$ $77.60_{\pm 0.25}$
TTM-RE (Gao et al., 2024) $78.13_{\pm 0.12}$ $78.05_{\pm 0.17}$ $79.95_{\pm 0.13}$ $78.20_{\pm 0.34}$
DREEAM (Ma et al., 2023) $79.42_{\pm 0.18}$ $78.36_{\pm 0.19}$ $80.20_{\pm 0.45}$ $78.56_{\pm 0.39}$
LLM-based (Li et al., 2024b)
LoRA FT LLaMA2-7B-Chat - - 53.0253.0253.02 52.7852.7852.78
LoRA FT LLaMA2-13B-Chat - - 52.4552.4552.45 52.1552.1552.15
LMRC LLaMA2-7B-Chat - - 72.9272.9272.92 72.3372.3372.33
LMRC LLaMA2-13B-Chat - - 74.6374.6374.63 74.0874.0874.08
LMRC LLaMA3-8B-Instruct - - 72.8072.8072.80 72.2172.2172.21
LMRC LLaMA3.1-8B-Instruct - - 72.9372.9372.93 72.3572.3572.35
Other GLiNER adaptation
GLiREL (Boylan et al., 2025) - - 54.1354.1354.13 53.2453.2453.24
GLiDRE (Ours) $77.76_{\pm 0.35}$ $76.70_{\pm 0.37}$ $77.83_{\pm 0.23}$ $76.80_{\pm 0.22}$
The results of our experiments presented in Table 1, demonstrate that GLiDRE achieves strong and competitive performance on the Re-DocRED benchmark. Our model attains a test F1 score of 77.8377.8377.83. This places GLiDRE in close competition with established DocRE models like ATLOP and KD-DocRE. While it does not surpass the current state-of-the-art methods like DREEAM and TTM-RE, it is important to note that these models incorporate additional mechanisms, such as evidence informations for DREEAM and the Turing Token Machine for TTM-RE that is designed to scale better than our approach with large datasets.
Critically, GLiDRE outperforms LLM-based methods. It surpasses the best-performing LMRC approach by over 3 F1 points, despite being a significantly smaller and more computationally efficient model. This highlights the strength of specialized encoder architectures for this task compared to fine-tuning general-purpose LLMs.
Furthermore, the comparison with GLiREL highlights the importance of our architectural adaptations for the document-level context. By designing representations specifically for relation extraction at the document level, GLiDRE achieves an improvement of over 23 F1 points.
3.3.2 Low-Resource Settings
In low-resource scenarios, we evaluate the efficiency or our model by comparing it against two powerful supervised baselines. We selected DREEAM, as it represents the state-of-the-art on the fully-supervised benchmark, and ATLOP, whose architecture and performances share similarities with our approach. This comparison is designed to determine the dataset size at which these conventional supervised methods can match the performance of our model and quantifying the advantages of our approach in data-scarce regimes.
Table 2: Few-shot F1 results for GLiDRE, ATLOP and DREEAM across various training set sizes ($N$). All models are trained on the same 5 subsets for each value of $N$. Best results per column are shown in bold.
Model N = 1 N = 5 N = 10 N = 50 N = 100 N = 500 N = 1000
GLiDRE $\mathbf{24.45}_{\pm 6.37}$ $\mathbf{33.71}_{\pm 6.11}$ $\mathbf{41.73}_{\pm 2.94}$ $\mathbf{54.54}_{\pm 1.39}$ $\mathbf{60.04}_{\pm 0.50}$ $69.26_{\pm 0.37}$ $\mathbf{72.09}_{\pm 0.31}$
ATLOP $4.32_{\pm 3.19}$ $18.76_{\pm 4.93}$ $29.48_{\pm 3.91}$ $50.36_{\pm 1.18}$ $57.17_{\pm 0.37}$ $68.91_{\pm 0.43}$ $71.70_{\pm 0.20}$
DREEAM $4.27_{\pm 3.50}$ $16.02_{\pm 7.46}$ $27.07_{\pm 5.81}$ $52.05_{\pm 2.00}$ $58.14_{\pm 0.86}$ $\mathbf{69.32}_{\pm 0.31}$ $71.67_{\pm 0.32}$
The results presented in Table 2 clearly demonstrate the superiority of GLiDRE in data-scarce environments. In the extremely low-data regime (N $\leq$ 100), our model establishes a lead over both ATLOP and the state-of-the-art model, DREEAM. For instance, with only 10 training samples, GLiDRE achieves an F1 score of 41.7341.7341.73, surpassing DREEAM by a margin of over 14 F1 points. This significant advantage persists up to N=100 samples.
As the amount of training data increases to 500 and 1000 samples, the performance gap narrows, and the baselines become more competitive. This indicates that while conventional supervised methods eventually catch up, our model’s architecture provides a distinct and substantial advantage when labeled data is minimal.
3.3.3 Zero-Shot Setting
The zero-shot setting for Document-Level Relation Extraction remains a largely underexplored research area. Given the scarcity of existing baselines, we establish a benchmark by evaluating the performance of much larger open-weight Large Language Models. We selected good performing models from the Open LLM Leaderboard (Fourrier et al., 2024), specifically employing the instruction-tuned versions of Qwen 2.5 72B (Team, 2024), Mistral Large 123B [^3] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama 3.3 70B [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and the model we used for synthetic pretraining generation, Mistral-Small 24B. Our prompting methodology is inspired by recent work in zero-shot Information Extraction (Yuan et al., 2023). We annotate entity mentions directly within the input text using special tags and to ensure structured and reliable outputs, we constrain the generation process to a strict (head, relation, tail) triplet format. This constrained decoding is implemented using regular expressions within the vLLM inference framework (Kwon et al., 2023), which facilitates parsing and mitigates the risk of hallucination. The detailed prompt template used for this task is provided in Figure 2. We use a fixed temperature of 0.
The results of our zero-shot evaluation, detailed in Table 3, highlight the efficiency and effectiveness of GLiDRE. We establish a strong benchmark by comparing our model against several state-of-the-art LLMs.
Our primary finding is that GLiDRE, a model with only a few hundred million parameters, achieves performance that is competitive with LLMs orders of magnitude larger. With a test F1 score of 17.3217.3217.32, GLiDRE outperforms Llama 3.3 70B and Mistral-Small 24B. This demonstrates that for specialized tasks like DocRE, a focused bi-encoder architecture can rival the zero-shot reasoning capabilities of massive general-purpose models. Furthermore, the results show a dramatic improvement when compared to the score of GPT-3.5 Turbo reported by Xue et al. (2024), showing the rapid recent progress of LLMs in zero-shot Information Extraction tasks.
Beyond performance, GLiDRE offers significant practical advantages. The LLM baselines entail substantial computational overhead, requiring vast memory footprints (e.g., over 300 GB for Mistral-Large and at least 4 A100-80GB) and significant energy consumption for inference. In contrast, GLiDRE operates with a memory footprint of under 10 GB. Moreover, LLMs often require complex engineering, including sophisticated prompting and constrained decoding, to produce structured, parsable outputs and mitigate hallucinations, while our model natively produces structured predictions.
The performance of GLiDRE shows some variability, which may come from a misalignment between their general-purpose pre-training corpora and the specific domain and relations of the Re-DocRED dataset. Adapting its pre-training by generating synthetic data with relation types aligned to the target dataset, as was done with GLiNER for Personally Identifiable Informations [^5] https://huggingface.co/urchade/gliner_multi_pii-v1, could improve results. However, this would likely reduce its ability to generalize to other DocRE datasets.
Table 3: Zero-shot results on the Re-DocRED dev and test sets. Results for GPT-3.5 Turbo are from Xue et al. (2024). All other LLMs are instruction-tuned versions.
Model Dev F1 Dev Ign F1 Test F1 Test Ign F1
Large Language Model Baselines
Qwen 2.5 72B 18.24 18.09 18.00 17.86
Llama 3.3 70B 15.67 15.57 15.81 15.71
Mistral-Large 123B 18.49 18.33 18.61 18.50
Mistral-Small 24B 14.39 14.27 14.38 14.27
GPT-3.5 Turbo - - 6.68 -
Proposed Model
GLiDRE 16.72 16.37 17.32 16.41
Figure 2: The prompt template used for zero-shot DocRE inference with LLMs. Entity mentions are marked with special tags to ground the model and the output format is strictly constrained to triplets to ensure parsability and reduce invalid generations.
3.4 Analysis
3.4.1 LogSumExp vs. Mean Pooling
LogSumExp (LSE) pooling serves as a smooth approximation of the max pooling operation. The operation is defined as:
$\text{LSE}(\mathbf{x_{1},...,x_{n}})=\log\left(\sum_{i=1}^{n}\exp(x_{i})\right)$
This pooling strategy was successfully applied to document-level relation extraction by Jia et al. (2019) and later adopted by ATLOP, where it showed slightly better performances over conventional mean pooling.
However, our empirical results, shown in Table 4, indicate a different outcome within our architecture. For GLiDRE, conventional mean pooling outperforms LSE by nearly a full F1 point on the test set. We attribute this discrepancy to architectural differences; unlike ATLOP’s bilinear classifier, GLiDRE employs a bi-encoder framework that compares relation and type embeddings. This structural divergence suggests that optimizations are not always directly transferable between models. Given that the original ATLOP paper reported only minor gains from LSE, our findings confirm that mean pooling is a more effective and robust choice for our model.
Table 4: Comparison of F1 scores on the test set using LogSumExp versus Mean pooling for entity mention aggregation.
Pooling Method Test F1 Test Ign F1
LogSumExp $76.86_{\pm 0.15}$ $75.72_{\pm 0.20}$
Mean $\mathbf{77.83}_{\pm 0.23}$ $\mathbf{76.80}_{\pm 0.22}$
3.4.2 Effect of Pre-training and Localized Context Pooling
We conduct an ablation study to isolate the individual contributions of our synthetic pre-training stage and the Localized Context Pooling (noted LOP in the ATLOP model) mechanism. Two variants of our model are evaluated: one finetuned without pre-training and another without the LOP module.
The results in Table 5 confirm that both components positively contribute to the final performance of GLiDRE. Removing the pre-training stage leads to the most significant performance decrease, with a drop of nearly 0.70.70.7 F1 points, underscoring the effectiveness of our synthetic data generation for model initialization. Disabling LOP results in a drop of 0.20.20.2 F1 points, which validates its role in refining relation representations. These findings justify the inclusion of both techniques in our final model architecture.
Table 5: Ablation study on the Test set. We report F1 scores after removing the pre-training stage and the Localized Context Pooling (LOP) module.
Configuration Test F1 Test Ign F1
GLiDRE $\mathbf{77.83}_{\pm 0.23}$ $\mathbf{76.80}_{\pm 0.22}$
w/o Pre-training $77.15_{\pm 0.42}$ $75.96_{\pm 0.49}$
w/o LOP $77.61_{\pm 0.09}$ $76.48_{\pm 0.11}$
3.4.3 Adapting the Adaptive Threshold Loss
ATLOP introduced an adaptive thresholding mechanism to learn a dynamic, per-relation decision boundary, thereby avoiding a suboptimal, fixed global threshold. This is achieved by adding a special "threshold" (TH) class to the classifier; a relation is predicted only if its logit surpasses that of the TH class.
Adapting this to GLiDRE is non-trivial due to our model’s bi-encoder architecture, which lacks a fixed classifier head. To replicate the mechanism, we introduce a small multi-layer perceptron (MLP) to predict the threshold logit. The MLP input is a concatenation of the candidate relation’s embedding and a global context vector, formed by averaging all possible relation type embeddings. The model is then trained using the original ATLOP loss function.
To save compute resources, this experiment is conducted on the model variant without pre-training. As shown in Table 6, this adaptation proved detrimental, degrading performance by approximately 1.41.41.4 F1 points. We hypothesize several reasons for this negative result: (1) our method of computing the threshold logit via a separate MLP is fundamentally different from ATLOP’s integrated classifier approach; (2) the ATLOP loss may be less effective at handling the severe class imbalance in DocRED compared to the Focal Loss used in our main model; (3) a dynamic threshold may be unnecessary for our model. We observed that the optimal global threshold for GLiDRE consistently converges near 0.5 and further tuning per-class thresholds on the development set did not improve test set performance.
Table 6: Comparison between standard training with Focal Loss and our adaptation of the ATLOP adaptive thresholding method. Experiments are conducted without the pre-training phase.
Configuration Test F1 Test Ign F1
Focal Loss $\mathbf{77.15}_{\pm 0.42}$ $\mathbf{75.96}_{\pm 0.49}$
Adaptive Threshold $75.74_{\pm 1.05}$ $74.84_{\pm 1.05}$
4 Conclusion
We present GLiDRE, a novel lightweight bi-encoder model for document-level relation extraction that reconceptualizes Document-level Relation Extraction as a direct representation matching problem. We show, through extensive experiments on the Re-DocRED benchmark, that GLiDRE achieves state-of-the-art few-shot performance and matches or surpasses much larger LLMs in zero-shot settings, all while operating with a fraction of their computational footprint.
GLiDRE not only rivals fully supervised baselines under low-resource regimes but also delivers highly structured predictions without complex prompting or constrained decoding. Its efficiency makes it a practical choice for real-world IE applications.
Future work will explore synthetic relation generation to further close remaining performance gaps and investigate dynamic thresholding methods tailored to the bi-encoder setup to enhance robustness across diverse relation schemas.
Limitations
Despite its efficiency and strong few-shot performance, GLiDRE faces several inherent limitations. First, the number of candidate relation pairs grows quadratically with the number of entities in a document, which can dramatically increase memory consumption for texts with high entity density and even lead to out-of-memory errors. Second, although the bi-encoder design can in principle accommodate longer contexts, it remains bound by the maximum sequence length of the underlying document encoder (e.g. 512 tokens for DeBERTa). Documents that exceed this limit must be truncated or segmented into chunks, potentially disrupting long-distance dependencies and harming performance.
Third, GLiDRE’s independent scoring of each entity-relation pair overlooks inter-label interactions. As a result, it can struggle to distinguish semantically similar relation types in the absence of a joint classification mechanism. Finally, our evaluation follows standard DocRE protocols by assuming gold entities and coreference chains are provided. Since GLiDRE does not perform named entity recognition or coreference resolution, its effectiveness in a fully end-to-end pipeline would depend on the accuracy of upstream modules, which may propagate errors and degrade overall performance.
Acknowledgments
This publication was made possible by the use of the FactoryIA supercomputer, financially supported by the Ile-De-France Regional Council. It also benefited from the support of the DataFIX project, financed by the French government under the France 2030 Programme and operated by Bpifrance.
Appendix A Appendix
A.1 Pre-training Data Generation
The synthetic data for the pre-training phase is generated using the prompt template detailed in Figure 3.
Figure 3: The prompt template used for pre-training data generation with Mistral-Small.
A.2 Zero-Shot Relation Extraction
To assess the generalization capabilities of our document-level model, we evaluated GLiDRE on the sentence-level zero-shot benchmarks, FewRel and Wiki-ZSL. These datasets mainly feature a single candidate entity pair per sentence, making the task closer to relation classification. Following standard protocols, we evaluate on splits where the set of $m$ test relations is disjoint from the relations seen during training.
The results are presented in Table 7. Despite being designed and pre-trained for the more complex document-level setting, GLiDRE demonstrates respectable performance. For a small number of unseen relations (m=5), our model is highly competitive, outperforming several strong baselines and GPT-4o.
However, as the number of unseen relations increases, GLiDRE’s performance degrades more rapidly than models specifically architected for sentence-level zero-shot relation classification, such as GLiREL and TMC-BERT. We attribute this to our model’s pre-training on multi-mention instances, which may make it less specialized for the sentence-level relation classification. Furthermore, its bi-encoder design also scores each relation on its own, making it harder to tell apart similar relation types.
$m$ Model Wiki-ZSL FewRel
P R F1 P R F1
5 RelationPrompt Chia et al. (2022) 70.66 83.75 76.63 90.15 88.50 89.30
DSP-ZRSC Lv et al. (2023) 94.10 77.10 84.80 93.40 92.50 92.90
ZSRE Tran et al. (2023) 94.50 96.48 95.46 96.36 96.68 96.51
MC-BERT Lan et al. (2023) 80.28 84.03 82.11 90.82 91.30 90.47
TMC-BERT Möller and Usbeck (2024) 90.11 87.89 88.92 93.94 93.30 93.62
GPT-4o 91.24 72.07 80.03 96.75 83.05 89.20
GLiREL Boylan et al. (2025) 89.41 80.67 83.28 96.84 93.41 94.20
GLiDRE 93.17 92.15 92.24 93.68 92.14 92.15
10 RelationPrompt Chia et al. (2022) 68.51 74.76 71.50 80.33 79.62 79.96
DSP-ZRSC Lv et al. (2023) 80.00 74.00 76.90 80.70 88.00 84.20
ZSRE Tran et al. (2023) 85.43 88.14 86.74 81.13 82.24 81.68
MC-BERT Lan et al. (2023) 72.81 73.96 73.38 86.57 85.27 85.92
TMC-BERT Möller and Usbeck (2024) 81.21 81.27 81.23 84.42 84.99 85.68
GPT-4o 77.62 66.14 68.35 84.07 58.00 66.20
GLiREL Boylan et al. (2025) 89.87 81.56 83.67 91.09 87.42 87.60
GLiDRE 73.98 73.11 70.89 86.16 82.92 81.74
15 RelationPrompt NG Chia et al. (2022) 54.45 29.43 37.45 66.49 40.05 49.38
DSP-ZRSC Lv et al. (2023) 77.50 64.40 70.40 82.90 78.10 80.40
ZSRE Tran et al. (2023) 64.68 65.01 65.30 66.44 69.29 67.82
MC-BERT Lan et al. (2023) 65.71 67.11 66.40 80.71 79.84 80.27
TMC-BERT Möller and Usbeck (2024) 73.62 74.07 73.77 82.11 79.93 81.00
GPT-4o 81.04 32.06 41.57 84.42 65.76 70.70
GLiREL Boylan et al. (2025) 79.44 74.81 73.91 88.14 84.69 84.48
GLiDRE 67.29 67.67 64.98 79.37 76.05 75.50
Table 7: Zero-shot performance comparison on the Wiki-ZSL and FewRel datasets for a varying number of unseen relations ($m$). Baseline results are reported from their respective original publications. GPT-4o results are from Boylan et al. (2025).