### v2

# GLiDRE: Generalist Lightweight model for Document-level Relation Extraction

Abstract
Relation Extraction (RE) is a fundamental task in Natural Language Processing, and its document-level variant poses significant challenges, due to complex interactions between entities across sentences. While supervised models have achieved strong results in fully resourced settings, their behavior with limited training data remains insufficiently studied. We introduce GLiDRE, a new compact model for document-level relation extraction, designed to work efficiently in both supervised and few-shot settings. Experiments in both low-resource supervised training and few-shot meta-learning benchmarks show that our approach outperforms existing methods in data-constrained scenarios, establishing a new state-of-the-art in few-shot document-level relation extraction. Our code will be publicly available [^1] https://github.com/robinarmingaud/glidre.
GLiDRE: Generalist Lightweight model for Document-level Relation Extraction
Robin Armingaud, Romaric Besançon Université Paris-Saclay, CEA, List, F-91120, Palaiseau, France {robin.armingaud, romaric.besancon}@cea.fr
1 Introduction
Document-level Relation Extraction (DocRE) is a challenging task in natural language processing that involves identifying relationships between entities across multiple sentences. In contrast, traditional sentence-level relation extraction, evaluated on supervised datasets like TACRED (Zhang et al., 2017) and SemEval-2010 Task 8 (Hendrickx et al., 2010) or few-shot benchmarks such as FewREL (Han et al., 2018) and Wiki-ZSL (Chen and Li, 2021), typically considers a single entity pair per instance, often without negative examples (i.e. sentences where no relation is present). These settings closely resemble relation classification rather than relation extraction. While recent works improve existing benchmarks, like FewRel 2.0 (Gao et al., 2019) which addresses the negative examples issue, they still operate under the constraint of classifying a fixed pair of entities.
DocRE presents a more realistic and complex scenario, especially relevant for real-world applications such as biomedical or industrial Information Extraction, where relationships often span sentence boundaries and the entity pairs of interest are not pre-identified. The complexity of DocRE is due to the quadratic growth of negative pairs with the number of entities in a document. Evaluation is commonly conducted on datasets like DocRED or Re-DocRED (Yao et al., 2019; Tan et al., 2022b).
Figure 1: Model architecture. GLiDRE uses a dual-encoder structure: a text encoder generates token embeddings, while a label encoder computes embeddings for relation types. An entity representation is formed by pooling the embeddings of its constituent tokens and mentions. For each entity pair, their representations are concatenated and processed by a feed-forward network before final scoring. The model can optionally incorporate a localized context pooling module (Zhou et al., 2021) to enrich the relation representation.
However, if recent works tackle sentence-level relation extraction in low-resource settings (Boylan et al., 2025; Li et al., 2024a; Lan et al., 2023), there has been little research on DocRE in similar settings. In zero-shot settings, LLMs show strong capabilities for tasks such as NER and RE (Sainz et al., 2024; Zhou et al., 2023; Wang et al., 2023; Wei et al., 2023), but their performance on zero-shot DocRE is still limited (Li et al., 2023, Xue et al., 2024). On the other hand, lightweight zero-shot models have also achieved great performances for NER, such as GLiNER (Zaratiana et al., 2024), which exploits similarities between entity type representations and textual span representations in the latent space. This approach allows better representations than LLMs due to its use of bidirectional encoders and it solves the scalability issues of autoregressive models, surpassing much larger LLMs in zero-shot settings. Moreover, as a zero-shot model, GLiNER is not restricted to a fixed set of relation labels and can generalize to arbitrary relation types at inference time. With the same inspiration, we offer the following contributions:
  • We propose GLiDRE, a new and efficient model for document-level relation extraction with zero or few annotated data, that leverages embedding similarities between relation types and entity pair representations.
  • We conduct extensive experiments on the Re-DocRED, FREDo and Re-FREDo benchmarks, demonstrating that GLiDRE achieves state-of-the-art results in challenging few-shot scenarios.
  • We demonstrate that smaller, specialized encoder models can offer an efficient alternative to larger LLMs for complex, document-level relation extraction and we present an up-to-date evaluation of these models on the Re-DocRED dataset in a zero-shot setting.
2 Related Work
Document-level RE methods
State-of-the-art supervised approaches to DocRE generally extend the model ATLOP (Zhou et al., 2021), which introduces Localized Context Pooling, using attention mechanisms to enhance relation representations, and an adaptive-thresholding loss that adjusts decision thresholds dynamically based on entity pairs. Subsequent work has extended ATLOP in various directions. For example, DREEAM (Ma et al., 2023) uses evidences to enhance performance, while KD-DocRE (Tan et al., 2022a) uses axial attention to model interdependencies among entity pairs, adopts an adaptive focal loss to address class imbalance, and integrates knowledge distillation. Other works enhance DocRE models with rule constraints, adding logical constraints to improve consistency, with techniques such as bidirectional rule mining Liu et al. (2023) or differentiable document-specific rule learning Zhang et al. (2025). Another trend leverages LLMs: for instance, LMRC Li et al. (2024b) combines a dedicated classifier to suggest candidate relations with the generative ability of a LLM for final decisions, significantly improving upon purely generative approaches.
Benchmarks for DocRE
Document-level Relation Extraction models are commonly evaluated on the benchmark DocRED Zhou et al. (2021), and its refined version Re-DocRED (Tan et al., 2022b), which reduces the issue of false negatives. These two datasets contain complex documents with an average of 19 entities and approximately 390 candidate pairs per text, for a total of 3,053 annotated Wikipedia articles in the training set. Beyond these benchmarks, other datasets have been proposed for specialized domains, such as HacRED (Cheng et al., 2021) for Chinese relation extraction and SciERC (Luan et al., 2018) for scientific literature. However, to the best of our knowledge, DocRED and Re-DocRED remain the only widely adopted English benchmarks that provide evidence annotations, an essential feature for fair comparison with evidence-aware models such as DREEAM.
Few-Shot Document-level Relation Extraction
Given the complexity of document-level relation extraction, the few-shot setting has received comparatively limited attention. A straightforward strategy involves training under data-constrained regimes, where only a small number of labeled samples are available. Moving beyond this paradigm, Popovic and Färber (2022) formulate Few-Shot Document-level Relation Extraction (FSDRE) as a meta-learning problem and introduce the FREDo dataset, constructed upon DocRED and SciERC. Models are evaluated on episodic tasks comprising support and query sets derived from DocRED in the original FREDo, and from Re-DocRED in its refined version, Re-FREDo (Meng et al., 2023). SciERC is employed as an out-of-domain test set to evaluate cross-domain generalization. Approaches evaluated on these benchmarks, such as RAPL (Meng et al., 2023) and TPN (Zhang and Kang, 2024), typically employ prototype-based learning to derive robust class representations from only a handful of examples.
3 Methodology
3.1 Document Level Relation Extraction
Let a document be represented as a sequence of tokens. Within this document, there is a set of $m$ pre-identified entities $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each entity $e_{i}$ is associated with one or more mentions in the text.
The objective of DocRE is to identify valid relational triplets. Each triplet is of the form $(e_{h},r,e_{t})$, where $e_{h},e_{t}\in E$ are the head and tail entities respectively ($h$ can be equal to $t$), and $r$ is the relation that holds between them. Based on this definition, DocRE is a multi-label classification task on every entity pair, where labels are the relation types.
3.2 GLiDRE architecture
GLiDRE architecture is inspired by the bi-encoder variant of GLiNER [^2] https://blog.knowledgator.com/meet-the-new-zero-shot-ner-architecture-30ffc2cb1ee0. The GLiDRE model, presented in Figure 1, uses two separate transformers, one for encoding the document and another for encoding relation labels, which offers several benefits: unlike the uni‑encoder version, the bi-encoder approach avoids concatenating labels and document tokens into a single input. Instead, we encode the document and labels separately. This eliminates the need for special separation tokens and overcomes encoder context length limits, permitting an effectively unlimited label set. Moreover, label embeddings can be precomputed, accelerating inference. However, the bi‑encoder architecture increases memory usage and lacks cross‑attention between labels. As a result, it may struggle to disambiguate semantically similar labels.
Label and word representations
Label embeddings are obtained by mean-pooling over the token representations constituting label names. The computed label embeddings are passed through a two-layer feedforward network if the output dimension of the label encoder differs from the configured latent space dimension of GLiDRE. For document tokens, we adopt the strategy used in standard NER models: for words split into subwords, we extract the representation of the first subword.
Relation representations
Given contextual embeddings $H\in\mathbb{R}^{L\times D}$ from the document encoder, where $L$ is the sequence length and $D$ is the hidden dimension of the document encoder, we construct a representation for each candidate relation defined by a pair of entities.
First, for each mention of an entity, the mention representation is computed by pooling the embeddings of its constituent words.
If an entity has multiple mentions, its final representation $h_{e}$ is derived by taking the mean of all its mention representations.
Finally, the representation $h_{r}$ for the relation between $e_{h}$ and $e_{t}$, is obtained by concatenating their respective entity representations and passing them through a two-layer feedforward network.
$h_{r}=\text{FFN}(h_{e_{h}}\otimes h_{e_{t}})$ (1)
where $\otimes$ denotes vector concatenation.
Relation Scoring
To determine whether a candidate entity pair $(e_{h},e_{t})$ instantiates a given relation type $t$, we measure the following matching score between the relation representation $h_{r}$ and the embedding of relation type $h_{t}$:
$s(e_{h},e_{t},t)=\sigma(h_{r}^{\top}h_{t})$ (2)
where $\sigma(x)=(1+e^{-x})^{-1}$ is the sigmoid function. A relation $t$ is assigned to the entity pair $(e_{h},e_{t})$ if
$s(e_{h},e_{t},t)>\tau$ (3)
where $\tau$ is a predefined decision threshold.
3.3 Loss Function
To address the issue of class imbalance within DocRE datasets, we use the Focal Loss function proposed by Lin et al. (2017) during training. The Focal Loss is a dynamically scaled cross-entropy loss designed to down-weight the contribution of well-classified examples and concentrating the training on hard examples. It is defined as:
$\text{FL}(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ (4)
where $p_{t}$ represents the model’s estimated probability for the ground-truth class. The focusing parameter $\gamma$ adjusts the rate at which easy examples are down-weighted and the weighting factor $\alpha_{t}$ balances the importance of positive and negative examples.
3.4 Localized Context Pooling
To create more refined representations, we also explore and adapt methods inspired by ATLOP Zhou et al. (2021) including Localized Context Pooling. Localized Context Pooling uses the attention scores from the encoder to focus on the most relevant parts of the document for a given entity pair.
Let $A\in\mathbb{R}^{L\times L}$ be the document-level attention matrix from the last encoder layer, averaged across all attention heads. For the head entity $e_{h}$ and tail entity $e_{t}$, we extract their corresponding attention vectors $A_{h}$ and $A_{t}$ by averaging the attention scores over their mention tokens. A joint attention vector $\alpha$ is computed via an element-wise product:
$\alpha=A_{h}\odot A_{t}$ (5)
This $\alpha$ is then normalized and used to compute a weighted sum of the contextual embeddings, creating a localized context vector $c_{\text{loc}}$:
$c_{\text{loc}}=\sum_{i=1}^{L}\frac{\alpha_{i}}{\sum_{j=1}^{L}\alpha_{j}}H_{i}$ (6)
Finally, the refined representations for the head $h^{\prime}_{e_{h}}$ and tail $h^{\prime}_{e_{t}}$ entities are created by concatenating their initial representations with the localized context vector and passing them through distinct two-layer feedforward network:
$h^{\prime}_{e_{h}}=\tanh(\text{FFN}_{h}(h_{e_{h}}\otimes c_{\text{loc}}))$ (7)
$h^{\prime}_{e_{t}}=\tanh(\text{FFN}_{t}(h_{e_{t}}\otimes c_{\text{loc}}))$ (8)
The final relation representation $h_{r}$ is then computed from $h^{\prime}_{e_{h}}$ and $h^{\prime}_{e_{t}}$ using Equation 1.
3.5 Pretraining data
For a better generalization, GLiDRE benefits from pretraining on a large-scale generalist annotated corpus. We build such a corpus using a semi-automated annotation methodology inspired by Zhou et al. (2023). Documents are first randomly sampled from FineWeb Penedo et al. (2024), a high-quality dataset built from filtered and deduplicated English Common Crawl archives.
We then prompt a LLM to generate annotations for both entities and the relations between them in a structured JSON format (we used the Mistral-Small-24B-Instruct-2501 model, the detailed prompt is provided in Appendix A.1). The raw outputs are filtered to remove instances containing malformed JSON or documents that exceed a predefined length threshold.
The resulting dataset contains 136,404 documents with highly diverse labels, featuring 76,497 unique relation types. The most frequent relation types include types such as IS_LOCATED_IN, WORKS_FOR, PART_OF and CONTAINS. We use this dataset to pretrain our model.
4 Experiments
4.1 Datasets
We conduct our experiments on the English benchmark Re-DocRED Tan et al. (2022b), a human-revised version of the DocRED benchmark that addresses its high false-negative rate, logical inconsistencies and coreference errors by re‑annotating all 4,053 documents in the original training and evaluation splits. Re-DocRED retains the same relation schema but substantially increases the number of annotated relation triples per document.
For evaluating performance in a formal few-shot meta-learning context, we additionally use the FReDo (Popovic and Färber, 2022) and Re-FReDo (Meng et al., 2023) benchmarks. These datasets re-purpose documents from DocRED, Re-DocRED, and SciERC into an episodic format. The evaluation is structured into two tasks: an in-domain task with approximately 15k episodes derived from DocRED, and a more challenging cross-domain task with 3k episodes derived from SciERC. Each episode consists of a small support set (1 or 3 documents) and a query set. To ensure a rigorous few-shot evaluation, the relation types in the training, development, and test splits are disjoint.
4.2 Evaluation Settings
Standard supervised settings
We evaluate the performance of our model on Re-DocRED across 3 distinct experimental settings, with a particular focus on low-data regimes where its advantages are most pronounced. First, to assess data efficiency generalization, we finetune different models on randomly sampled subsets of the training data containing 1, 5, 10, 50, 100, 500, or 1000 documents. Second, in a fully supervised setting, we finetune on the entire training set and compare against strong baselines. Finally, we examine zero-shot performance by comparing against larger LLMs without any task-specific fine-tuning, highlighting the competitiveness of the model in scenarios with little or no labeled data.
Episodic meta-learning settings
To evaluate our model in the FSDRE setting on the FReDo and Re-FReDo benchmarks, we adopted a two-stage fine-tuning protocol for a fair comparison with prototype-based methods. First, we perform an initial alignment phase by training the model for 500 steps on the full training set, mainly to align label representations (e.g., capitalization consistency). Subsequently, for each episode, we further finetune the model for 20 epochs on the provided 1-shot or 3-shot support set before evaluating its performance on the corresponding query set.
Evaluation Metrics
We use the standard metrics for the Re-DocRED dataset (Tan et al., 2022b): F1 is the micro-averaged F1 score on the relation triples, Ign_F1 is the F1 score ignoring the triples in the test set that appear also in the training set.
Implementation Details
All experiments are conducted on a single NVIDIA H100 GPU with a batch size of 16. For the pretraining phase, the model is trained for 50,000 steps. For finetuning, we train for 10,000 steps. We use two learning rates $1\times 10^{-5}$ for the model encoders and $1\times 10^{-4}$ for all other layers. For the low-resource and fully supervised experiments, we report the average and standard deviation over 5 runs with different seeds. The model checkpoint yielding the best performance on the development set and a decision threshold of 0.5 are used for the evaluation on the test set. We use two English models: DebertaV3 Large He et al. (2021) for document encoder and BGE-Large V1.5 Xiao et al. (2023) for label encoder, similar to the bi-encoder version of GLiNER. The final model comprises approximately 800 million parameters. The pretraining phase requires approximately 24 hours, while finetuning on the ReDocRED dataset takes 3.5 hours. Additionally, evaluating our model on the Re-FReDo benchmark is computationally intensive due to the episodic protocol, which requires repeated fine-tuning and model weight resets. This process amounts to approximately 35 hours of computation for the in-domain task and 5 hours for the cross-domain task.
4.3 Results
We present in this section the main results of the evaluation of GLiDRE in several settings. An additional detailed analysis of some parameters of the model such as the effect of pretraining, pooling strategies and adaptive thresholding are presented in Appendix A.3.
4.3.1 Low-Resource Settings
In low-resource scenarios, we evaluate the efficiency of our model by comparing it against two strong supervised baselines:
DREEAM (Ma et al., 2023) represents state-of-the-art on the fully-supervised benchmark and relies on guiding the model attention with evidence information as supervisory signals in a memory‐efficient manner and can employs a self‐training strategy to learn evidence retrieval without explicit annotations.
ATLOP (Zhou et al., 2021) formulates document-level relation extraction as a semantic‐segmentation task, introducing localized context pooling to capture entity‐focused local contexts and an adaptive thresholding loss to learn dynamic decision thresholds. It is more comparable to our approach since it does not requires additional evidence information.
This comparison is designed to determine the dataset size at which these conventional supervised methods can match the performance of our model and quantifying the advantages of our approach in data-scarce regimes.
Model N = 1 N = 5 N = 10 N = 50 N = 100 N = 500 N = 1000
ATLOP $4.32_{\pm 3.19}$ $18.76_{\pm 4.93}$ $29.48_{\pm 3.91}$ $50.36_{\pm 1.18}$ $57.17_{\pm 0.37}$ $68.91_{\pm 0.43}$ $71.70_{\pm 0.20}$
DREEAM $4.27_{\pm 3.50}$ $16.02_{\pm 7.46}$ $27.07_{\pm 5.81}$ $52.05_{\pm 2.00}$ $58.14_{\pm 0.86}$ $\mathbf{69.32}_{\pm 0.31}$ $71.67_{\pm 0.32}$
GLiDRE (Ours) $\mathbf{24.45}_{\pm 6.37}$ $\mathbf{33.71}_{\pm 6.11}$ $\mathbf{41.73}_{\pm 2.94}$ $\mathbf{54.54}_{\pm 1.39}$ $\mathbf{60.04}_{\pm 0.50}$ $69.26_{\pm 0.37}$ $\mathbf{72.09}_{\pm 0.31}$
Table 1: Low-resource supervised setting: micro-averaged F1 results for GLiDRE, ATLOP and DREEAM across various training set sizes (NN). All models are trained on the same 5 subsets for each value of $N$. Best results per column are shown in bold.
The results presented in Table 1 clearly demonstrate the superiority of GLiDRE in data-scarce environments. In the extremely low-data regime (N $\leq$ 100), our model establishes a lead over both ATLOP and the state-of-the-art model, DREEAM. For instance, with only 10 training samples, GLiDRE achieves an F1 score of $41.73$, surpassing DREEAM by a margin of over 14 F1 points. This significant advantage persists up to N=100 samples, the performance gap narrowing only when the amount of training data increases to 500 and 1000 samples.
4.3.2 Few-Shot Document-level Relation Extraction
Model FREDo ReFREDo
In-Domain Cross-Domain In-Domain Cross-Domain
1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$ 1-Doc $F_{1}$ 3-Doc $F_{1}$
DL-Base 0.60 0.89 1.76 1.98 1.38 1.84 1.76 1.98
DL-MNAV 7.05 8.42 0.84 0.48 12.97 12.43 1.12 2.28
$\textnormal{DL-MNAV}_{SIE}$ 7.06 6.77 1.77 2.51 13.37 12.00 1.39 2.92
$\textnormal{DL-MNAV}_{SIE+SBN}$ 1.71 2.79 2.85 3.72 4.59 5.43 2.84 3.86
RAPL 8.75 10.67 3.33 5.35 15.20 16.35 3.51 5.48
TPN 9.12 8.64 3.98 4.48 15.54 15.73 4.72 5.02
GLiDRE (Ours) 13.00 15.71 10.97 11.00 26.36 28.77 10.60 10.32
Table 2: Few-shot episodic setting: results on FREDo and ReFREDo benchmarks. Reported scores are macro-averaged across relation types and obtained from Meng et al. (2023) and Zhang and Kang (2024). Best results per column are shown in bold.
For the FSDRE evaluation, we follow the episodic protocols defined by FREDo and Re-FREDo. We compare GLiDRE to three methods specifically developed for episodic few-shot DocRE:
DL-MNAV (Popovic and Färber, 2022) adapts sentence-level few-shot relation extraction MNAV Sabo et al. (2021) to documents by pooling mention representations and explicitly modeling NOTA (none-of-the-above) via learned NOTA vectors, adaptive-threshold loss, and support-based NOTA sampling at inference.
RAPL (Meng et al., 2023) refines prototypes with instance-level aggregation and relation-weighted contrastive learning, and constructs task-specific NOTA prototypes.
TPN (Zhang and Kang, 2024) improves cross-domain transfer with a hybrid encoder, transferable NOTA prototype learning, and a calibration module to mitigate NOTA bias.
As shown in Table 2, GLiDRE establishes a new state-of-the-art on both benchmarks by a substantial margin across all settings. The advantage of our model is particularly pronounced in the challenging cross-domain setting, where the task is to generalize to relations from the SciERC dataset, highlighting the effectiveness of our synthetic data pretraining strategy for FSDRE and domain transfer.
Table 3: Fully supervised setting: results on the Re-DocRED dataset. We compare GLiDRE with strong DocRE models, LLM-based methods and GLiREL. F1 scores are reported on both the development and test sets. Results are taken from the original papers, except for DocRE models, which are reported following Gao et al. (2024).
Model Dev F1 Dev Ign F1 Test F1 Test Ign F1
DocRE Models
ATLOP (Zhou et al., 2021) $76.15_{\pm 0.23}$ $5.88_{\pm 0.23} 77$$7.81_{\pm 0.71}$ $76.13_{\pm 0.28}$
KD-DocRE (Tan et al., 2022a) $7.88_{\pm 0.42} 77$$7.12_{\pm 0.49}$ $8.28_{\pm 0.72} 77$$7.60_{\pm 0.25}$
TTM-RE (Gao et al., 2024) $78.13_{\pm 0.12}$ $78.05_{\pm 0.17}$ $79.95_{\pm 0.13}$ $78.20_{\pm 0.34}$
DREEAM (Ma et al., 2023) $79.42_{\pm 0.18}$ $78.36_{\pm 0.19}$ $80.20_{\pm 0.45}$ $78.56_{\pm 0.39}$
LLM-based (Li et al., 2024b)
LMRC LLaMA2-13B-Chat - - $74.63$ $74.08$
GLiNER-inspired model
GLiREL (Boylan et al., 2025) - - $54.13$ $53.24$
GLiDRE (Ours) $77.76_{\pm 0.35}$ $6.70_{\pm 0.37} 77$$7.83_{\pm 0.23}$ $76.80_{\pm 0.22}$
4.3.3 Fully supervised setting
Additionally, we evaluate the scalability of our proposed model, GLiDRE, by comparing it against several strong baselines on the Re-DocRED dataset under a fully supervised protocol. Specifically, in addition to DREEAM and ATLOP, we also include the following baselines:
KD-DocRE (Tan et al., 2022a) leverages an axial attention module to model entity-pairs interdependency across sentences, applies an adaptive focal loss to mitigate class imbalance and uses knowledge distillation to incorporate distantly supervised data.
TTM-RE (Gao et al., 2024) introduces a Token Turing Machine memory module that augments document representations with external memory tokens, paired with a noise‐robust loss tailored for positive–unlabeled setting.
Additionally, we benchmark against GLiREL (Boylan et al., 2025), a sentence-level relation extraction adaptation of GLiNER architecture. It is important to note that GLiREL was originally conceived for sentence-level relation extraction and operates with a shorter context length. For document-level predictions, it relies on aggregating relations based on gold-standard coreference information.
We also report recent methods that leverage LLMs, comparing with the framework introduced by Li et al. (2024b), which involves fine-tuning Llama models (Touvron et al., 2023) using Low-Rank Adaptation (Hu et al., 2022) and enhancing the model performance by employing a classifier to exhibit potential relations and guide the finetuned LLM. We only report their best result on this dataset.
The results presented in Table 3, demonstrate that GLiDRE achieves competitive performance on the Re-DocRED benchmark. Notably, this performance is attained despite the model not being specifically designed for large-scale datasets, highlighting its versatility and suitability across diverse scenarios. Our model attains a test F1 score of $77.83$. This places GLiDRE in close competition with established DocRE models like ATLOP and KD-DocRE. While it does not surpass the current state-of-the-art methods like DREEAM and TTM-RE, it is important to note that these models incorporate additional mechanisms, such as evidence information for DREEAM and the Turing Token Machine for TTM-RE that is designed to scale better than our approach with large datasets.
Critically, GLiDRE outperforms LLM-based methods. It surpasses the best-performing LMRC approach by over 3 F1 points, despite being a significantly smaller and more computationally efficient model. This highlights the strength of specialized encoder architectures for this task compared to fine-tuning general-purpose LLMs.
Furthermore, the comparison with GLiREL highlights the importance of our architectural adaptations for the document-level context. By designing representations specifically for relation extraction at the document level, GLiDRE achieves an improvement of over 23 F1 points.
4.3.4 Zero-Shot Setting
Table 4: Zero-shot setting: results on the Re-DocRED dev and test sets. Results for GPT-3.5 Turbo are from Xue et al. (2024). All other LLMs are instruction-tuned versions.
Model Dev F1 Dev Ign F1 Test F1 Test Ign F1
Large Language Model Baselines
Qwen 2.5 72B 18.24 18.09 18.00 17.86
Llama 3.3 70B 15.67 15.57 15.81 15.71
Mistral-Large 123B 18.49 18.33 18.61 18.50
Mistral-Small 24B 14.39 14.27 14.38 14.27
GPT-3.5 Turbo - - 6.68 -
Proposed Model
GLiDRE 16.72 16.37 17.32 16.41
The zero-shot setting for Document-Level Relation Extraction remains a largely underexplored research area. Given the scarcity of existing baselines, we establish a benchmark by evaluating the performance of much larger open-weight Large Language Models. We compare against recent high-performing models, specifically employing the instruction-tuned versions of Qwen 2.5 72B (Team, 2024), Mistral Large 123B [^3] https://huggingface.co/mistralai/Mistral-Large-Instruct-2411, Llama 3.3 70B [^4] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct and the model we used for pretraining data generation, Mistral-Small 24B. Our prompting methodology is inspired by recent work in zero-shot Information Extraction (Yuan et al., 2023). We annotate entity mentions directly within the input text using special tags and, in order to ensure structured and reliable outputs, we constrain the generation process to a strict (head, relation, tail) triplet format. This constrained decoding is implemented using regular expressions within the vLLM inference framework (Kwon et al., 2023), which facilitates parsing and mitigates the risk of hallucination. The detailed prompt template used for this task is provided in Appendix A.2. We use a fixed temperature of 0.
The results of our zero-shot evaluation, detailed in Table 4, highlight the efficiency and effectiveness of GLiDRE. We establish a strong benchmark by comparing our model against several state-of-the-art LLMs.
Our primary finding is that GLiDRE, a model with only a few hundred million parameters, achieves performance that is competitive with LLMs orders of magnitude larger. With a test F1 score of $17.32$, GLiDRE outperforms Llama 3.3 70B and Mistral-Small 24B (which was used to generate the pretraining data) demonstrating the effectiveness to train on the pretraining corpus, as opposed to directly leveraging Mistral-Small 24B to annotate DocRED. This demonstrates that for specialized tasks like DocRE, a focused bi-encoder architecture can rival the zero-shot reasoning capabilities of massive general-purpose models. Furthermore, the results show a dramatic improvement when compared to the score of GPT-3.5 Turbo reported by Xue et al. (2024), showing the rapid recent progress of LLMs in zero-shot Information Extraction tasks.
Beyond performance, GLiDRE offers significant practical advantages. The inference process on the Re-DocRED test set (500 documents) with GLiDRE (0.8B parameters) requires less than 10GB of VRAM and completes in approximately 100 seconds on a single NVIDIA A100-80GB GPU. In contrast, inference with Mistral-Large (123B parameters) necessitates a distributed setup of at least 4x A100-80GB GPUs to accommodate its memory footprint of over 300GB. This process takes around 600 seconds and a total cost of approximately 2400 GPU-seconds. This breakdown demonstrates that GLiDRE is not only smaller but over 20 times more cost-effective in terms of total compute required for inference. Moreover, using LLMs for specific tasks often require complex engineering, including sophisticated prompting and constrained decoding, to produce structured, parsable outputs and mitigate hallucinations, while our model natively produces structured predictions.
The performance of GLiDRE shows some variability, which may come from a misalignment between their general-purpose pretraining corpora and the specific domain and relations of the Re-DocRED dataset. Adapting its pretraining by generating synthetic data with relation types aligned to the target dataset, as was done with GLiNER for Personally Identifiable Informations [^5] https://huggingface.co/urchade/gliner_multi_pii-v1, could improve results. However, this would likely reduce its ability to generalize to other DocRE datasets.
5 Conclusion
We present GLiDRE, a novel lightweight bi‑encoder model for document‑level relation extraction that reconceptualizes Document-level Relation Extraction as a direct representation matching problem. We show, through extensive experiments on the Re‑DocRED, FREDo and Re-FREDo benchmarks, that GLiDRE achieves state‑of‑the‑art few‑shot performance and matches or surpasses much larger LLMs in zero‑shot settings, all while operating with a fraction of their computational footprint.
GLiDRE not only rivals fully supervised baselines under low‑resource regimes but also delivers highly structured predictions without complex prompting or constrained decoding. Its efficiency makes it a practical choice for real‑world IE applications.
Future work will explore synthetic relation generation to further close remaining performance gaps and investigate dynamic thresholding methods tailored to the bi‑encoder setup to enhance robustness across diverse relation schemas.
Limitations
Despite its efficiency and strong few‑shot performance, GLiDRE faces several inherent limitations.
The number of candidate relation pairs grows quadratically with the number of entities in a document, which can dramatically increase memory consumption for texts with high entity density and even lead to out‑of‑memory errors.
Although the bi‑encoder design can in principle accommodate longer contexts, it remains bound by the maximum sequence length of the underlying document encoder (e.g. 512 tokens for DeBERTa). Other approaches we compare against, such as DREEAM, ATLOP, KD-DocRE, and TTM-RE, have the same limitation. Documents that exceed this limit must be truncated or segmented into chunks, potentially disrupting long‑distance dependencies and harming performance.
The independent scoring of each entity‐relation pair in GLiDRE overlooks inter‑label interactions. As a result, the model can struggle to distinguish semantically similar relation types in the absence of a joint classification mechanism.
Our evaluation follows standard DocRE protocols by assuming gold entities and coreference chains are provided. Since GLiDRE does not perform named entity recognition or coreference resolution, its effectiveness in a fully end‑to‑end pipeline would depend on the accuracy of upstream modules, which may propagate errors and degrade overall performance. This limitation is shared by other existing baselines and is particularly pronounced for DREEAM, which additionally relies on evidence annotations.
The pretraining of GLiDRE relies on the synthetic annotation of texts with a LLM, which is a source of risk for the propagation of the inherent biases of the LLM. This risk is mitigated by the use of clear annotation guidelines and could be further limited by combining multiple prompts and models and having human reviewers audit a subset of the annotations for bias and correctness.
Acknowledgments
This publication was made possible by the use of the FactoryIA supercomputer, financially supported by the Ile-De-France Regional Council. It also benefited from the support of the DataFIX project, financed by the French government under the France 2030 Programme and operated by Bpifrance.
Appendix A Appendix
A.1 Pretraining Data Generation
The synthetic data for the pretraining phase is generated using the prompt template detailed in Figure 2. An example of generated data is provided in Figure 3
Figure 2: The prompt template used for pretraining data generation with Mistral-Small.
Figure 3: Example of generated data.
A.2 Zero-shot DocRE inference with LLMs
Figure 4: The prompt template used for zero-shot DocRE inference with LLMs. Entity mentions are marked with special tags to ground the model and the output format is strictly constrained to triplets to ensure parsability and reduce invalid generations.
A.3 Analysis
A.3.1 LogSumExp vs. Mean Pooling
LogSumExp (LSE) pooling serves as a smooth approximation of the max pooling operation. The operation is defined as:
$\text{LSE}(\mathbf{x_{1},...,x_{n}})=\log\left(\sum_{i=1}^{n}\exp(x_{i})\right)$
This pooling strategy was successfully applied to document-level relation extraction by Jia et al. (2019) and later adopted by ATLOP, where it showed slightly better performances over conventional mean pooling.
However, our empirical results, shown in Table 5, indicate a different outcome within our architecture. For GLiDRE, conventional mean pooling outperforms LSE by nearly a full F1 point on the test set. We attribute this discrepancy to architectural differences; unlike the bilinear classifier of ATLOP, GLiDRE employs a bi-encoder framework that compares relation and type embeddings. This structural divergence suggests that optimizations are not always directly transferable between models. Given that the original ATLOP paper reported only minor gains from LSE, our findings confirm that mean pooling is a more effective and robust choice for our model.
Table 5: Comparison of F1 scores on the test set using LogSumExp versus Mean pooling for entity mention aggregation.
Pooling Method Test F1 Test Ign F1
LogSumExp $76.86_{\pm 0.15}$ $75.72_{\pm 0.20}$
Mean $\mathbf{77.83}_{\pm 0.23}$ $\mathbf{76.80}_{\pm 0.22}$
A.3.2 Effect of pretraining and Localized Context Pooling
We conduct an ablation study to isolate the individual contributions of our synthetic pretraining stage and the Localized Context Pooling (noted LOP in the ATLOP model) mechanism. Two variants of our model are evaluated: one finetuned without pretraining and another without the LOP module.
The results in Table 6 confirm that both components positively contribute to the final performance of GLiDRE. Removing the pretraining stage leads to the most significant performance decrease, with a drop of nearly $0.7$ F1 points, underscoring the effectiveness of our synthetic data generation for model initialization. Disabling LOP results in a drop of $0.2$ F1 points, which validates its role in refining relation representations. These findings justify the inclusion of both techniques in our final model architecture.
Table 6: Ablation study on the Test set. We report F1 scores after removing the pretraining stage and the Localized Context Pooling (LOP) module.
Configuration Test F1 Test Ign F1
GLiDRE $\mathbf{77.83}_{\pm 0.23}$ $\mathbf{76.80}_{\pm 0.22}$
w/o pretraining $77.15_{\pm 0.42}$ $75.96_{\pm 0.49}$
w/o LOP $77.61_{\pm 0.09}$ $76.48_{\pm 0.11}$
A.3.3 Adapting the Adaptive Threshold Loss
ATLOP introduced an adaptive thresholding mechanism to learn a dynamic, per-relation decision boundary, thereby avoiding a suboptimal, fixed global threshold. This is achieved by adding a special "threshold" (TH) class to the classifier; a relation is predicted only if its logit surpasses that of the TH class.
Adapting this to GLiDRE is non-trivial due to our bi-encoder architecture, which lacks a fixed classifier head. To replicate the mechanism, we introduce a small multi-layer perceptron (MLP) to predict the threshold logit. The MLP input is a concatenation of the candidate relation embedding and a global context vector, formed by averaging all possible relation type embeddings. The model is then trained using the original ATLOP loss function.
To save compute resources, this experiment is conducted on the model variant without pretraining. As shown in Table 7, this adaptation proved detrimental, degrading performance by approximately $1.4$ F1 points. We hypothesize several reasons for this negative result: (1) our method of computing the threshold logit via a separate MLP is fundamentally different from ATLOP’s integrated classifier approach; (2) the ATLOP loss may be less effective at handling the severe class imbalance in DocRED compared to the Focal Loss used in our main model; (3) a dynamic threshold may be unnecessary for our model. We observed that the optimal global threshold for GLiDRE consistently converges near 0.5 and further tuning per-class thresholds on the development set did not improve test set performance.
Table 7: Comparison between standard training with Focal Loss and our adaptation of the ATLOP adaptive thresholding method. Experiments are conducted without the pretraining phase.
Configuration Test F1 Test Ign F1
Focal Loss $\mathbf{77.15}_{\pm 0.42}$ $\mathbf{75.96}_{\pm 0.49}$
Adaptive Threshold $75.74_{\pm 1.05}$ $74.84_{\pm 1.05}$
A.4 Zero-Shot Sentence-level Relation Extraction
To assess the generalization capabilities of our document-level model, we evaluated GLiDRE on the sentence-level zero-shot benchmarks, FewRel and Wiki-ZSL. These datasets mainly feature a single candidate entity pair per sentence, making the task closer to relation classification. Following standard protocols, we evaluate on splits where the set of $m$ test relations is disjoint from the relations seen during training.
The results are presented in Table 8. Despite being designed and pre-trained for the more complex document-level setting, GLiDRE demonstrates respectable performance. For a small number of unseen relations (m=5), our model is highly competitive, outperforming several strong baselines and GPT-4o.
However, as the number of unseen relations increases, GLiDRE performance degrades more rapidly than models specifically architected for sentence-level zero-shot relation classification, such as GLiREL and TMC-BERT. We attribute this to our model pretraining on multi-mention instances, which may make it less specialized for the sentence-level relation classification. Furthermore, its bi‑encoder design also scores each relation on its own, making it harder to tell apart similar relation types.
$m$ Model Wiki-ZSL FewRel
P R F1 P R F1
5 RelationPrompt Chia et al. (2022) 70.66 83.75 76.63 90.15 88.50 89.30
DSP-ZRSC Lv et al. (2023) 94.10 77.10 84.80 93.40 92.50 92.90
ZSRE Tran et al. (2023) 94.50 96.48 95.46 96.36 96.68 96.51
MC-BERT Lan et al. (2023) 80.28 84.03 82.11 90.82 91.30 90.47
TMC-BERT Möller and Usbeck (2024) 90.11 87.89 88.92 93.94 93.30 93.62
GPT-4o 91.24 72.07 80.03 96.75 83.05 89.20
GLiREL Boylan et al. (2025) 89.41 80.67 83.28 96.84 93.41 94.20
GLiDRE 93.17 92.15 92.24 93.68 92.14 92.15
10 RelationPrompt Chia et al. (2022) 68.51 74.76 71.50 80.33 79.62 79.96
DSP-ZRSC Lv et al. (2023) 80.00 74.00 76.90 80.70 88.00 84.20
ZSRE Tran et al. (2023) 85.43 88.14 86.74 81.13 82.24 81.68
MC-BERT Lan et al. (2023) 72.81 73.96 73.38 86.57 85.27 85.92
TMC-BERT Möller and Usbeck (2024) 81.21 81.27 81.23 84.42 84.99 85.68
GPT-4o 77.62 66.14 68.35 84.07 58.00 66.20
GLiREL Boylan et al. (2025) 89.87 81.56 83.67 91.09 87.42 87.60
GLiDRE 73.98 73.11 70.89 86.16 82.92 81.74
15 RelationPrompt NG Chia et al. (2022) 54.45 29.43 37.45 66.49 40.05 49.38
DSP-ZRSC Lv et al. (2023) 77.50 64.40 70.40 82.90 78.10 80.40
ZSRE Tran et al. (2023) 64.68 65.01 65.30 66.44 69.29 67.82
MC-BERT Lan et al. (2023) 65.71 67.11 66.40 80.71 79.84 80.27
TMC-BERT Möller and Usbeck (2024) 73.62 74.07 73.77 82.11 79.93 81.00
GPT-4o 81.04 32.06 41.57 84.42 65.76 70.70
GLiREL Boylan et al. (2025) 79.44 74.81 73.91 88.14 84.69 84.48
GLiDRE 67.29 67.67 64.98 79.37 76.05 75.50
Table 8: Zero-shot performance comparison on the Wiki-ZSL and FewRel datasets for a varying number of unseen relations ($m$). Baseline results are reported from their respective original publications. GPT-4o results are from Boylan et al. (2025).